size Ion 0.01 0.05 0.001 0.005 0.1 a,pm) Charge +/-1 Activity coefficient (y) H+ 900 0.967 0.933 0.914 0.86 0.83 |(C,HS2СHCO, (С,Н, ) (02N)3CH20-(CH,)NH+, CH30gH4CO2- 800 0.966 0.931 0.912 0.82 0.85 700 0.965 0.930 0.909 0.845 0.81 u+,CgHgco,, носдH, со,, CIC, H,Cо,, CsH,CH,со2, сH,-снсH,со,-, (CH3)2CHCH2CO2-, (CH3CH2)4N+, (C3H7)2NH2+ 600 0.907 0.965 0.929 0.835 0.80 C,CHCO ,, Cl, cco (CH,CH2),NH+, (с,н, NH,+ |Na+, CdCi+, cIo,, 10;, нсо,, н,РОд, HSO,, H,As0, Co(NH3)4(NO2)2, CH3C02, CICH2C02, (CH3)4N*, (CH3CH2)2NH2+, H2NCH2CO2 |+н,NCH,CO, H, (Cн,), NH. сH, сH,NH3+ 500 0.964 0.928 0.904 0.83 0.79 450 0.928 0.902 0.964 0.82 0.775 400 0.964 0.927 0.77 0.901 0.815 |он-, F-, SCN-, оCN-, HS-, CIO,, CIO, Bro,, 10,,, MпO4, HCо,, ttocH H2citrate, CH3NH3+, (CH3)2NH2+ 0.926 0.76 350 0.964 0.900 0.81 K+, C, Br, I, CN-, NO2, NO3 Rb+, Cs+, NH+, TI+, Ag+ 300 0.964 0.899 0.925 0.805 0.755 0.964 0.924 0.898 0.80 250 0.75 Activity coefficient (y) Charge +/-2 Mq2+, Be2+ 800 0.872 0.69 0.755 0.52 0.45 |сH2(CH,CH,CO, )z (CH,CH,CH,CO,)2 Ca2+, Cu2+, Zn2+, Sn2+, Mn2+, Fet2, Ni2+, Co+2, CH4(CO2)2 Н.С(CH,CO, )2, (CH,CH,CO,) Sr2+, Ba2+, Cd2+, Hg2+, S2-, S2042, WO42, H2C(CO2")2, (CH2CO22! (CнОнсо, )2 700 0.872 0.755 0.685 0.50 0.425 600 0.870 0.749 0.675 0.485 0.405 500 0.744 0.67 0.868 0.465 0.38 450 0.867 0.742 0.665 0.455 0.37 Pb2+, co32-, so32-, Mo042-, Co(NH3)5c2+, Fe(CN)5NO2-, c2042, Hcitrate2- Hg2+, SO2, S2032-, s2062-, s20g2-, Seo2-, Cro2, HPO,2- 0.867 0.740 400 0.660 0.445 0.355 Charge +/-3 A13+, Fe3+, Cr3+, Sc3+, Y3+, In3+, lanthanidesa Activity coefficient (y) 0.738 0.445 900 0.54 0.245 0.18 citrate3- 500 0.728 0.405 0.51 0.18 0.115 PO 3, Fe(CN)3, Cr(NH3)3+, Co(NH3)3+, Co(NH3)sH203+ 400 0.725 0.505 0.395 0.16 0.095 Charge+/- 4 Activity coefficient (y) Th4+, Zr4+, Ce4+, Sn4+ 1100 0.588 0.35 0.255 0.10 0.065 Fe(CN)4 500 0.57 0.20 0.31 0.048 0.021 a (3athanidec 1 3
Determine the pH of a solution containing 0.030 M N2OH and 0.025 M KI neglecting activities. pH Determine the pH of the same solution including activities. Activity coefficients at various ionic strengths can be found in this table pH =
Answer #1

Guuen data, Craouc O.030M FO 030 ue mow thut i - log (oH-) -log(o 030) POH =-6.s2) POH =152 PH 14-POH = 14-152 PH =12,4S we k

-osi oOSS + (3s0 00ss 30S -O 1196 -O 10 tto cue mow that POHog Co)804 -l0g Co023) Po C63) 163 P=14-POH 14-1-63 Py 12.37 cs S